Shopping at Super Sale Days, Martha buys her children 3 shirts and 2 pairs of pants for $85.50. She returns during the sale and buys 4 more shirts and 3 more pairs of pants for $123. What is the sale price of the shirts and pants?
Answer by FrankM(1040)

3S+2P=85.50
4S+3P=123
Let's multiply the first equation by 3 and the second by 2
9S+6P=256.50
8S+6P=246.00
Subtract the second equation from the first.
S=10.50
and put that back into first equation
3(10.50)+2P=85.50
31.50+2P=85.50
2P=54
P=27
Shirts are $10.50 Pants $27.00